Solve the following system by the substitution method.
x+3y=2
-x+y=1

Answer by edjones(8007)

x+3y=2
-x+y=1
.
y=x+1
x+3(x+1)=2 substitute x+1 for y
x+3x+3=2
4x=-1
x=-1/4
y=3/4
.
Ed